# RL-100: Performant Robotic Manipulation with Real-World Reinforcement Learning

### v2

@@ Abstract
Real-world robotic manipulation in homes and factories demands reliability, efficiency, and robustness that approach or surpass skilled human operators. We present RL-100, a real-world reinforcement learning training framework built on diffusion visuomotor policies trained by supervised learning. RL-100 introduces a three-stage pipeline. First, imitation learning leverages human priors. Second, iterative offline reinforcement learning uses an Offline Policy Evaluation procedure, abbreviated OPE, to gate PPO-style updates that are applied in the denoising process for conservative and reliable improvement. Third, online reinforcement learning eliminates residual failure modes. An additional lightweight consistency distillation head compresses the multi-step sampling process in diffusion into a single-step policy, enabling high-frequency control with an order-of-magnitude reduction in latency while preserving task performance. The framework is task-, embodiment-, and representation-agnostic and supports both 3D point clouds and 2D RGB inputs, a variety of robot platforms, and both single-step and action-chunk policies. We evaluate RL-100 on seven real-robot tasks spanning dynamic rigid-body control, such as Push-T and Agile Bowling, fluids and granular pouring, deformable cloth folding, precise dexterous unscrewing, and multi-stage orange juicing. RL-100 attains 100% success across evaluated trials for a total of 900 out of 900 episodes, including up to 250 out of 250 consecutive trials on one task. The method achieves near-human teleoperation or better time efficiency and demonstrates multi-hour robustness with uninterrupted operation lasting up to two hours. Policies generalize zero-shot to novel dynamics with an average 90.0% success, adapt in a few-shot manner to significant task variations (86.7% after 1–3 hours of additional training), and remain robust under aggressive human perturbations (about 95% success). Notably, our juicing robot served random customers continuously for about 7 hours without failure when zero-shot deployed in a shopping mall. These results suggest a practical path to deployment-ready robot learning by starting from human priors, aligning training objectives with human-grounded metrics, and reliably extending performance beyond human demonstrations. For more results and videos, please visit our project website: https://lei-kun.github.io/RL-100/.
@@ 1 Introduction
Dexterous robotic manipulation stands as an iconic challenge in robotics (Cui and Trinkle, 2021; Luo et al., 2025). Real-world deployment beyond laboratories requires human-level reliability, efficiency, and robustness. Recent learning-based advances across generative diffusion policies (Chi et al., 2023; Ze et al., 2024), diffusion-based robot foundation models (Black et al., 2024a; Intelligence et al., 2025), sim-to-real RL (Yuan et al., 2025; Lin et al., 2025), and real-world RL (Luo et al., 2025, 2024) have narrowed this gap, demonstrating human-like manipulation proficiency. In particular, generative policies and foundation models are trained or fine-tuned on high-quality, human-collected real-robot datasets at varying scales, providing strong human priors and enabling robots to acquire the efficient strategies used by skilled tele-operators. However, high-quality real-robot data remain scarce: teleoperation incurs perceptual and control latency and favors slow, conservative motions (Guo et al., 2025). Moreover, large-scale collection depends on skilled operators and is labor-intensive and expensive. As a result, state–action coverage is limited, undermining generalization and reliability. Consequently, this supervised paradigm is constrained by an imitation ceiling: under purely supervised objectives, performance is effectively bounded by demonstrator skill and inherits human inefficiencies, biases, and occasional errors.
@@
During policy rollouts, a human operator gives sparse success signals when needed, and the controller follows conservative operating limits. We use a unified policy-gradient objective across both iterative offline and online phases to fine-tune the diffusion sampler’s short-horizon denoising schedule (Ho et al., 2020a). This alignment yields stable updates across phases and strong fine-tuning sample efficiency. In addition, we interleave a lightweight distillation loss that compresses the $K$-step diffusion policy into a one-step consistency (Song et al., 2023) policy for deployment, reducing inference latency while maintaining or improving efficiency and robustness.
Moreover, our framework is task- and embodiment-agnostic. We evaluate RL-100 across simulated tasks and on a real-world suite of seven manipulation tasks as illustrated in Fig. 1 and summarized in Tab. 1: Dynamic Push-T, Agile Bowling, Pouring, Soft-towel Folding, Dynamic Unscrewing, and Orange Juicing. Orange Juicing comprises two subtasks, placing and removal, which are trained and evaluated separately but reported as one task family. The suite includes rigid-body dynamics, deformable objects, fluids, and precision assembly, and the framework is deployed across multiple embodiments. For these tasks and embodiments, we select a specific control mode per task: a single-step control mode is used when a fast closed-loop reaction is necessary; action-chunk-based control is preferred for coordination-heavy or high precision tasks where smoothing mitigates jitter and limits error compounding (Zhao et al., 2023). Both regimes share the same diffusion backbone; only the action heads differ.
@@ 2.1 Reinforcement Learning
Our post-training follows an offline-to-online paradigm. Following LEI et al. (2024), we employ a proximal policy optimization (PPO)-style objective (Schulman et al., 2017; Zhuang et al., 2023) to unify both stages. The core learning objective incorporates importance sampling for proximal policy updates:
@@
  • Offline: Implicit Q Learning (IQL)-style (Kostrikov et al., 2022) value functions: $A^{\text{off}}(s,a)=Q(s,a)-V(s)$.
  • Online: Generalized Advantage Estimation (GAE) (Schulman et al., 2016): $A^{\text{on}}(s,a)=\text{GAE}(R_{t},V)$ to balance variance and bias.
@@ 2.2 Diffusion models
We overload the subscript $t$ from timestep in MDP to the step index in diffusion process in the following two subsections. Diffusion models (Ho et al., 2020b) learn to reverse a noising process that gradually corrupts clean data $x_{0}\!\in\!\mathbb{R}^{d}$ into Gaussian noise to reconstruct the original clean data distribution. Given a schedule $\{\alpha_{t}\}_{t=1}^{T}$ with $\alpha_{t}=1-\beta_{t}$, a sample $x_{0}$ drawn from the clean distribution, the forward noising process follows a closed form:
@@ 2.3 DDIM sampling with stochastic form
Denoising Diffusion Implicit Models (DDIM) (Ho et al., 2020a) provide a family of samplers that interpolate between deterministic and stochastic generation. Given a learned denoiser $\varepsilon_{\theta}$, the predicted clean sample at time $t$ is commonly written as
@@ 2.4 Consistency Models
Consistency models (Song et al., 2023) learn a single-step mapping from noisy inputs at arbitrary noise levels to clean data. Denote the consistency model by $C_{\theta}(x^{\tau},\tau)$ where the superscript indicates the diffusion index (per the notation above). Given a frozen diffusion teacher $\Psi_{\varphi}$ (for instance a $K$-step DDIM teacher that follows the same subsampled schedule $\{\tau_{k}\}$), consistency distillation minimizes the squared regression objective
@@ 2.5 Diffusion policy and RL fine-tuning
Diffusion Policy (Chi et al., 2023) performs diffusion over robot actions conditioning on observations. Given an observation $o$, we roll out along the $K$-step subsampled schedule $\{\tau_{K}>\cdots>\tau_{1}\}$:
@@
Now we embed the $K$-step diffusion process as a sub-MDP into a single step in robotic manipulation MDP (Ren et al., 2024). Each denoising step can be viewed as sampling a cleaner noisy action $a^{\tau_{k-1}}$ from the Gaussian sub-policy in (7a). We therefore model the process as a $K$-step sub-MDP with:
@@
The log-likelihood defined in (7b) then computes the density of sub-policies, enabling end-to-end optimization of task rewards with respect to the denoising sub-policies via policy-gradient updates (e.g., PPO (Schulman et al., 2017)).
@@ 3.1 Imitation Learning
For RGB input, we use pretrained ResNet/ViT backbones; for point clouds input, we adapt the 3D encoder of DP3 (Ze et al., 2024) with reconstruction regularization for stability during RL fine-tuning. Visual embeddings are projected and concatenated with proprioceptive features to form $c_{t}$. All encoders are trained end-to-end with Eq. (14).
@@ 3.2.1 Offline RL
where the critics $(Q_{\psi},V_{\psi})$ are pre-trained on $\mathcal{D}$ following IQL (Kostrikov et al., 2022). This yields a candidate policy $\pi$ from $\pi_{i}$ by several epochs of gradient updates on $\mathcal{D}$.
@@
We use AM-Q (LEI et al., 2024) for offline policy evaluation (OPE) without further interaction with the environment to compare the candidate with current behavior policy:
@@ 4.1 Reinforcement Learning with Generative Diffusion Models
The integration of generative diffusion models with RL represents a paradigm shift in policy representation and optimization. Building upon foundational work in diffusion models (Ho et al., 2020a; Song et al., 2020) and flow matching (Lipman et al., 2023), recent advances have demonstrated the power of these generative frameworks in capturing complex, multimodal action distributions inherent in RL problems. Diffusion Q-Learning (DQL) (Wang et al., 2023) pioneered this integration by replacing traditional Gaussian policies with conditional diffusion models in offline reinforcement learning, addressing fundamental limitations of parametric policies in modeling multimodal behaviors. This approach has evolved through multiple directions: weighted regression methods (Kang et al., 2023; Lu et al., 2023; Ding et al., 2024) train diffusion policies through importance-weighted objectives to maximize learned Q-functions; reparameterization gradient approaches (Psenka et al., 2024; He et al., 2023; Ding and Jin, 2023) utilize gradient-based optimization despite temporal backpropagation challenges; and sampling-based methods (Chen et al., 2024) provide effective but computationally expensive solutions. More recently, consistency-based extensions(Li et al., 2024) have further generalized diffusion and consistency policies to visual RL.
Recent works have also explored using RL to directly optimize diffusion models beyond traditional likelihood-based training. Black et al. (2024b) demonstrate fine-tuning diffusion models with RL to maximize arbitrary reward functions, while Fan et al. (2023) apply similar techniques specifically to text-to-image generation using human feedback. Ren et al. (2024) introduce policy gradient methods tailored for diffusion-based policies, enabling effective online optimization while maintaining the expressiveness benefits of diffusion models.
To address computational bottlenecks inherent in diffusion models, Park et al. (2025) present Flow Q-Learning (FQL), which leverages flow-matching policies to model complex action distributions while avoiding recursive backpropagation through denoising processes. By training an independent single-step policy that matches the flow model’s output, FQL achieves computational efficiency without sacrificing expressiveness, demonstrating state-of-the-art performance on offline RL benchmarks with significant computational advantages, particularly in offline-to-online fine-tuning settings. Similarly, One-Step Flow Q-Learning (OFQL) (Nguyen and Yoo, 2025) extends this framework for even faster single-step action generation.
@@ 4.2 Generative Diffusion Models in Robotics
The application of diffusion models to robotics has yielded transformative advances in visuomotor control, trajectory planning, and real-world deployment. Diffusion Policy (Chi et al., 2023) demonstrated breakthrough performance in visuomotor control by modeling action distributions as conditional diffusion processes, excelling at handling multimodal behaviors through visual conditioning and receding horizon control. This approach has inspired extensions including DiffClone (Mani et al., 2024) and Diffusion Model-Augmented Behavioral Cloning (Chen et al., 2023), which further improve upon traditional behavioral cloning by leveraging the expressiveness of diffusion models. The integration with pretrained visual representations, including R3M (Nair et al., 2022), VC-1 (Majumdar et al., 2023), and MoCo (He et al., 2020), has proven particularly effective for generalization across visual variations. Recent work on 3D Diffusion Policy (Ze et al., 2024) extends these capabilities to point cloud observations, while FlowPolicy (Zhang et al., 2025a) introduces consistency flow matching for robust 3D manipulation tasks. Lu et al. (2025) propose a triply-hierarchical diffusion policy that decomposes complex visuomotor tasks into multiple levels of abstraction, improving both learning efficiency and generalization.
Beyond direct policy learning, diffusion models have revolutionized trajectory planning in robotics. Janner et al. (2022) and Ajay et al. (2022) reformulate planning as conditional generation, producing complete state-action trajectories conditioned on rewards and constraints, excelling in long-horizon tasks requiring complex coordination. Hegde et al. (2025) introduces an alternative paradigm by generating policy parameters rather than trajectories in latent spaces, offering computational advantages. Real-world deployment challenges have been addressed through methods like One-Step Diffusion Policy (Wang et al., 2025), which uses distillation to achieve real-time performance suitable for robotic control. Successful applications now span deformable object manipulation with PinchBot (Bartsch et al., 2025), multi-task learning, and navigation scenarios. The availability of large-scale datasets like BridgeData V2 (Walke et al., 2023) and foundation models such as $\pi_{0}$ (Black et al., 2024a) enables broader generalization across robotic platforms and tasks, accelerating the transition from laboratory demonstrations to practical robotic systems.
@@ 4.3 Offline-to-Online Reinforcement Learning
The transition from offline pretraining to online fine-tuning presents unique challenges in managing distribution shift and preventing catastrophic forgetting while enabling continuous improvement. Conservative Q-Learning (CQL) (Kumar et al., 2020) established the foundational framework for safe offline RL through pessimistic value estimation, preventing overestimation for out-of-distribution actions. Advantage-Weighted Regression (AWR) (Peng et al., 2019) provides a scalable framework that has influenced numerous subsequent works, though its restriction to Gaussian policies limits expressiveness for complex behavioral patterns. Calibrated Q-Learning (Cal-QL) (Nakamoto et al., 2023) addresses initialization challenges by ensuring conservative Q-values are appropriately scaled for effective online fine-tuning, providing crucial insights for successful offline-to-online transfer. Uni-O4 (LEI et al., 2024) directly applies the PPO (Schulman et al., 2017) objective to unify offline and online learning, eliminating the need for extra regularization.
Recent advances focus on efficiently leveraging both offline and online data through hybrid strategies. RLPD (Ball et al., 2023) achieves sample-efficient online learning by mixing offline and online experiences, demonstrating that careful data mixture strategies can accelerate learning significantly. Nair et al. (2020) established fundamental frameworks for combining offline pretraining with online fine-tuning, showing that hybrid approaches achieve superior sample efficiency compared to pure online or offline methods. A paradigm shift in offline-to-online adaptation is represented by Wagenmaker et al. (2025), which introduces DSRL (Diffusion Steering with Reinforcement Learning) that operates RL entirely in the latent noise space of pretrained diffusion policies rather than modifying base policy weights.

### v3

@@ Abstract
Real-world robotic manipulation in homes and factories demands reliability, efficiency, and robustness that approach or surpass skilled human operators. We present RL-100, a real-world reinforcement-learning framework built on diffusion visuomotor policies. RL-100 unifies imitation and reinforcement learning under a single PPO-style objective applied within the denoising process, yielding conservative, stable improvements across offline and online stages. To meet deployment latency, a lightweight consistency distillation compresses multi-step diffusion into a one-step controller for high-frequency control. The framework is task-, embodiment-, and representation-agnostic, and supports both single-action and action-chunking control. We evaluate RL-100 on seven diverse real-robot tasks—from dynamic pushing and agile bowling to pouring, cloth folding, unscrewing, and multi-stage juicing. RL-100 attains 100% success across evaluated trials for a total of 900 out of 900 episodes, including up to 250 out of 250 consecutive trials on one task. It matches or surpasses expert teleoperators in time-to-completion. Without retraining, a single policy attains approximately 90% zero-shot success under environmental and dynamics shifts, adapts in a few-shot regime to significant task variations (86.7%), and remains robust to aggressive human perturbations (about 95%). Notably, our juicing robot served random customers continuously for about seven hours without failure when zero-shot deployed in a shopping mall. These results suggest a practical path to deployment-ready robot learning by starting from human priors, aligning training objectives with human-grounded metrics, and reliably extending performance beyond human demonstrations. For more results and videos, please visit our project website: https://lei-kun.github.io/RL-100/.
@@ 1 Introduction
Dexterous robotic manipulation stands as an iconic challenge in robotics (cuisr; luosr). Real-world deployment beyond laboratories requires human-level reliability, efficiency, and robustness. Recent learning-based advances across generative diffusion policies (dp; dp3), diffusion-based robot foundation models (pi0; pi0.5), sim-to-real RL (zc1; toru), and real-world RL (luosr; luo2024serl) have narrowed this gap, demonstrating human-like manipulation proficiency. In particular, generative policies and foundation models are trained or fine-tuned on high-quality, human-collected real-robot datasets at varying scales, providing strong human priors and enabling robots to acquire the efficient strategies used by skilled tele-operators. However, high-quality real-robot data remain scarce: teleoperation incurs perceptual and control latency and favors slow, conservative motions (glx). Moreover, large-scale collection depends on skilled operators and is labor-intensive and expensive. As a result, state–action coverage is limited, undermining generalization and reliability. Consequently, this supervised paradigm is constrained by an imitation ceiling: under purely supervised objectives, performance is effectively bounded by demonstrator skill and inherits human inefficiencies, biases, and occasional errors.
@@
During policy rollouts, a human operator gives sparse success signals when needed, and the controller follows conservative operating limits. We use a unified policy-gradient objective across both iterative offline and online phases to fine-tune the diffusion sampler’s short-horizon denoising schedule (ho2020denoising). This alignment yields stable updates across phases and strong fine-tuning sample efficiency. In addition, we interleave a lightweight distillation loss that compresses the $K$-step diffusion policy into a one-step consistency (cm) policy for deployment, reducing inference latency while maintaining or improving efficiency and robustness.
Moreover, our framework is task- and embodiment-agnostic. We evaluate RL-100 across simulated tasks and on a real-world suite of seven manipulation tasks as illustrated in Fig. 1 and summarized in Tab. 1: Dynamic Push-T, Agile Bowling, Pouring, Soft-towel Folding, Dynamic Unscrewing, and Orange Juicing. Orange Juicing comprises two subtasks, placing and removal, which are trained and evaluated separately but reported as one task family. The suite includes rigid-body dynamics, deformable objects, fluids, and precision assembly, and the framework is deployed across multiple embodiments. For these tasks and embodiments, we select a specific control mode per task: a single-step control mode is used when a fast closed-loop reaction is necessary; action-chunk-based control is preferred for coordination-heavy or high precision tasks where smoothing mitigates jitter and limits error compounding (aloha). Both regimes share the same diffusion backbone; only the action heads differ.
@@ 2.1 Reinforcement Learning
Our post-training follows an offline-to-online paradigm. Following lei2024unio, we employ a proximal policy optimization (PPO)-style objective (ppo; bppo) to unify both stages. The core learning objective incorporates importance sampling for proximal policy updates:
@@
  • Offline: Implicit Q Learning (IQL)-style (iql) value functions: $A^{\text{off}}(s,a)=Q(s,a)-V(s)$.
  • Online: Generalized Advantage Estimation (GAE) (gae): $A^{\text{on}}(s,a)=\text{GAE}(R_{t},V)$ to balance variance and bias.
@@ 2.2 Diffusion models
We overload the subscript $t$ from timestep in MDP to the step index in diffusion process in the following two subsections. Diffusion models (ddpm) learn to reverse a noising process that gradually corrupts clean data $x_{0}\!\in\!\mathbb{R}^{d}$ into Gaussian noise to reconstruct the original clean data distribution. Given a schedule $\{\alpha_{t}\}_{t=1}^{T}$ with $\alpha_{t}=1-\beta_{t}$, a sample $x_{0}$ drawn from the clean distribution, the forward noising process follows a closed form:
@@ 2.3 DDIM sampling with stochastic form
Denoising Diffusion Implicit Models (DDIM) (ho2020denoising) provide a family of samplers that interpolate between deterministic and stochastic generation. Given a learned denoiser $\varepsilon_{\theta}$, the predicted clean sample at time $t$ is commonly written as
@@ 2.4 Consistency Models
Consistency models (cm) learn a single-step mapping from noisy inputs at arbitrary noise levels to clean data. Denote the consistency model by $C_{\theta}(x^{\tau},\tau)$ where the superscript indicates the diffusion index (per the notation above). Given a frozen diffusion teacher $\Psi_{\varphi}$ (for instance a $K$-step DDIM teacher that follows the same subsampled schedule $\{\tau_{k}\}$), consistency distillation minimizes the squared regression objective
@@ 2.5 Diffusion policy and RL fine-tuning
Diffusion Policy (dp) performs diffusion over robot actions conditioning on observations. Given an observation $o$, we roll out along the $K$-step subsampled schedule $\{\tau_{K}>\cdots>\tau_{1}\}$:
@@
Now we embed the $K$-step diffusion process as a sub-MDP into a single step in robotic manipulation MDP (ren2024diffusion). Each denoising step can be viewed as sampling a cleaner noisy action $a^{\tau_{k-1}}$ from the Gaussian sub-policy in (7a). We therefore model the process as a $K$-step sub-MDP with:
@@
The log-likelihood defined in (7b) then computes the density of sub-policies, enabling end-to-end optimization of task rewards with respect to the denoising sub-policies via policy-gradient updates (e.g., PPO (ppo)).
@@ 3.1 Imitation Learning
For RGB input, we use pretrained ResNet/ViT backbones; for point clouds input, we adapt the 3D encoder of DP3 (dp3) with reconstruction regularization for stability during RL fine-tuning. Visual embeddings are projected and concatenated with proprioceptive features to form $c_{t}$. All encoders are trained end-to-end with Eq. (14).
@@ 3.2.1 Offline RL
where the critics $(Q_{\psi},V_{\psi})$ are pre-trained on $\mathcal{D}$ following IQL (iql). This yields a candidate policy $\pi$ from $\pi_{i}$ by several epochs of gradient updates on $\mathcal{D}$.
@@
We use AM-Q (lei2024unio) for offline policy evaluation (OPE) without further interaction with the environment to compare the candidate with current behavior policy:
@@ 4.1 Reinforcement Learning with Generative Diffusion Models
The integration of generative diffusion models with RL represents a paradigm shift in policy representation and optimization. Building upon foundational work in diffusion models (ho2020denoising; song2020score) and flow matching (lipman2022flow), recent advances have demonstrated the power of these generative frameworks in capturing complex, multimodal action distributions inherent in RL problems. Diffusion Q-Learning (DQL) (wang2022diffusion) pioneered this integration by replacing traditional Gaussian policies with conditional diffusion models in offline reinforcement learning, addressing fundamental limitations of parametric policies in modeling multimodal behaviors. This approach has evolved through multiple directions: weighted regression methods (kang2023efficient; lu2023contrastive; ding2024diffusion) train diffusion policies through importance-weighted objectives to maximize learned Q-functions; reparameterization gradient approaches (psenka2024qscore; he2023diffusion; ding2024consistency) utilize gradient-based optimization despite temporal backpropagation challenges; and sampling-based methods (chen2023score) provide effective but computationally expensive solutions. More recently, consistency-based extensions(li2024cp3er) have further generalized diffusion and consistency policies to visual RL.
Recent works have also explored using RL to directly optimize diffusion models beyond traditional likelihood-based training. black2023training demonstrate fine-tuning diffusion models with RL to maximize arbitrary reward functions, while fan2024dpok apply similar techniques specifically to text-to-image generation using human feedback. ren2024diffusion introduce policy gradient methods tailored for diffusion-based policies, enabling effective online optimization while maintaining the expressiveness benefits of diffusion models.
To address computational bottlenecks inherent in diffusion models, park2025flow present Flow Q-Learning (FQL), which leverages flow-matching policies to model complex action distributions while avoiding recursive backpropagation through denoising processes. By training an independent single-step policy that matches the flow model’s output, FQL achieves computational efficiency without sacrificing expressiveness, demonstrating state-of-the-art performance on offline RL benchmarks with significant computational advantages, particularly in offline-to-online fine-tuning settings. Similarly, One-Step Flow Q-Learning (OFQL) (nguyen2025revisiting) extends this framework for even faster single-step action generation.
@@ 4.2 Generative Diffusion Models in Robotics
The application of diffusion models to robotics has yielded transformative advances in visuomotor control, trajectory planning, and real-world deployment. Diffusion Policy (dp) demonstrated breakthrough performance in visuomotor control by modeling action distributions as conditional diffusion processes, excelling at handling multimodal behaviors through visual conditioning and receding horizon control. This approach has inspired extensions including DiffClone (sabatelli2024diffclone) and Diffusion Model-Augmented Behavioral Cloning (wang2023diffusion), which further improve upon traditional behavioral cloning by leveraging the expressiveness of diffusion models. The integration with pretrained visual representations, including R3M (nair2022r3m), VC-1 (majumdar2023we), and MoCo (he2020momentum), has proven particularly effective for generalization across visual variations. Recent work on 3D Diffusion Policy (dp3) extends these capabilities to point cloud observations, while FlowPolicy (chen2024flowpolicy) introduces consistency flow matching for robust 3D manipulation tasks. h3dp propose a triply-hierarchical diffusion policy that decomposes complex visuomotor tasks into multiple levels of abstraction, improving both learning efficiency and generalization.
Beyond direct policy learning, diffusion models have revolutionized trajectory planning in robotics. janner2022planning and ajay2023conditional reformulate planning as conditional generation, producing complete state-action trajectories conditioned on rewards and constraints, excelling in long-horizon tasks requiring complex coordination. shang2024latent introduces an alternative paradigm by generating policy parameters rather than trajectories in latent spaces, offering computational advantages. Real-world deployment challenges have been addressed through methods like One-Step Diffusion Policy (wang2024onestep), which uses distillation to achieve real-time performance suitable for robotic control. Successful applications now span deformable object manipulation with PinchBot (liu2024pinchbot), multi-task learning, and navigation scenarios. The availability of large-scale datasets like BridgeData V2 (walke2023bridgedata) and foundation models such as $\pi_{0}$ (pi0) enables broader generalization across robotic platforms and tasks, accelerating the transition from laboratory demonstrations to practical robotic systems.
@@ 4.3 Offline-to-Online Reinforcement Learning
The transition from offline pretraining to online fine-tuning presents unique challenges in managing distribution shift and preventing catastrophic forgetting while enabling continuous improvement. Conservative Q-Learning (CQL) (kumar2020conservative) established the foundational framework for safe offline RL through pessimistic value estimation, preventing overestimation for out-of-distribution actions. Advantage-Weighted Regression (AWR) (peng2019advantage) provides a scalable framework that has influenced numerous subsequent works, though its restriction to Gaussian policies limits expressiveness for complex behavioral patterns. Calibrated Q-Learning (Cal-QL) (nakamoto2023calql) addresses initialization challenges by ensuring conservative Q-values are appropriately scaled for effective online fine-tuning, providing crucial insights for successful offline-to-online transfer. Uni-O4 (lei2024unio) directly applies the PPO (ppo) objective to unify offline and online learning, eliminating the need for extra regularization.
Recent advances focus on efficiently leveraging both offline and online data through hybrid strategies. RLPD (ball2023efficient) achieves sample-efficient online learning by mixing offline and online experiences, demonstrating that careful data mixture strategies can accelerate learning significantly. nair2020awac established fundamental frameworks for combining offline pretraining with online fine-tuning, showing that hybrid approaches achieve superior sample efficiency compared to pure online or offline methods. A paradigm shift in offline-to-online adaptation is represented by wagenmaker2025steering, which introduces DSRL (Diffusion Steering with Reinforcement Learning) that operates RL entirely in the latent noise space of pretrained diffusion policies rather than modifying base policy weights.